If you have infinite energy can you release infinite power forever by just releasing half of the remains energy each second?
1
Expert's answer
2019-06-11T10:03:26-0400

Infinity is not a number. That's why it does not obey the rules of algebra.

For example:


"\infty+\infty=\infty"

"\infty+1=\infty"

We cannot operate with infinite things in that manner: one infinity of this plus half infinity of that.